OpenStudy (anonymous):

Which of the following equations describe the line shown below? (1,10) and (-3,2) Check all that apply. A. y - 3 = 2(x - 2) B. y - 2 = 2(x - 3) C. y - 10 = 2(x - 1) D. y - 2 = 2(x + 3) E. y - 1 = 2(x - 10) F. y + 3 = 2(x - 2)

OpenStudy (anonymous):

the coordinates (1,10) and (-3,2) were from the graph.

OpenStudy (anonymous):

can you find your slope?

OpenStudy (acxbox22):

find slope then substitute one of the points in point slope equation

Nnesha (nnesha):

bec you answer choices and all answer choices have same slope so...:D:P but you can still find for more practice

OpenStudy (anonymous):

So it would be all of them?

Nnesha (nnesha):

nope lol slope is same not the y-intercept do you know point slope formula ??

Nnesha (nnesha):

if you don't know what is point slope equation then ask to @amanda7414 or @@acxbox22 they will help you best of luck@!!

Nnesha (nnesha):

nope slope should be 2 bec all answer choices have slope =2 here m is slope

OpenStudy (triciaal):

@Tallan you used the wrong numbers

OpenStudy (anonymous):

Oh!

OpenStudy (acxbox22):

you have m (slope) and now sub in the x and y values from any point into that equation above by any point i mean the 2 points given in the problem the y and x have to be of the same point - you cant have an x from one point and y from another

Nnesha (nnesha):

like i said yesterday use this formula \[m =\frac{ y _{2}-y _{1} }{ x _{2} -x _{1} }\ so in order pair (x1 , y1) (x2 , y2)

OpenStudy (triciaal):

using (x1, y1) = (1,10) and (x2, y2) = (-3, 2) y2-y1/x2-x1 m = (2 -10)/-3 -1)= then use this slope (x, y) and either of the points

Nnesha (nnesha):

 here it is